OpenStudy (ilikecookies):

what is a converse to this statement? : If x+11=5, then x=-6

OpenStudy (anonymous):

The converse of : if p then q is : if q then p

OpenStudy (anonymous):

The converse of "If it is cloudy, then it is raining." is "If it is raining, then it is cloudy."

OpenStudy (ilikecookies):

oooh is it x=-6 if x+11=5

OpenStudy (ilikecookies):

and i have another one if ur down to answer it

OpenStudy (anonymous):

no keep the if....then.... structure if x-6 then x+11=5

OpenStudy (anonymous):

i mean x = -6

OpenStudy (ilikecookies):

ohhhhhhhhhhhhhhhhhh

OpenStudy (ilikecookies):

but then you keep the x+11=5 the same??

OpenStudy (anonymous):

you don't change anything in the "if {A} then {B} "sentence when you do the converse. only switch the A and B.

OpenStudy (ilikecookies):

oh okay i get it do u have time for another?

OpenStudy (ilikecookies):

guess not